# Gradient-Direction-Aware Density Control for 3D Gaussian Splatting

### v1

@@ Abstract
The emergence of 3D Gaussian Splatting (3DGS) has significantly advanced novel view synthesis through explicit scene representation, enabling real-time photorealistic rendering. However, existing approaches manifest two critical limitations in complex scenarios: (1) Over-reconstruction occurs when persistent large Gaussians cannot meet adaptive splitting thresholds during density control. This is exacerbated by conflicting gradient directions that prevent effective splitting of these Gaussians; (2) Over-densification of Gaussians occurs in regions with aligned gradient aggregation, leading to redundant component proliferation. This redundancy significantly increases memory overhead due to unnecessary data retention. We present Gradient-Direction-Aware Gaussian Splatting (GDAGS), a gradient-direction-aware adaptive density control framework to address these challenges. Our key innovations: the gradient coherence ratio (GCR), computed through normalized gradient vector norms, which explicitly discriminates Gaussians with concordant versus conflicting gradient directions; and a nonlinear dynamic weighting mechanism leverages the GCR to enable gradient-direction-aware density control. Specifically, GDAGS prioritizes conflicting-gradient Gaussians during splitting operations to enhance geometric details while suppressing redundant concordant-direction Gaussians. Conversely, in cloning processes, GDAGS promotes concordant-direction Gaussian densification for structural completion while preventing conflicting-direction Gaussian overpopulation. Comprehensive evaluations across diverse real-world benchmarks demonstrate that GDAGS achieves superior rendering quality while effectively mitigating over-reconstruction, suppressing over-densification, and constructing compact scene representations with 50% reduced memory consumption through optimized Gaussians utilization. Our code can be accessed in the code repository https://github.com/zzcqz/GDAGS.
@@ 1 Introduction
Figure 1: Visualization of rendering performance and efficiency, with rendering frame rate and scene Gaussian count displayed in the upper left corner. The green box represents the optimized area, and the red box represents the poorly optimized area. Both GDAGS and AbsGS effectively densify and eliminate large Gaussians in over-reconstructed areas to restore fine details. However, AbsGS has too many over-densification Gaussians, which only contribute slightly to rendering quality. In contrast, GDAGS has a more compact and reasonable spatial distribution and only requires half of the Gaussians in Pixel-GS.
@@
Novel View Synthesis (NVS) constitutes a foundational challenge in computer graphics and vision, aiming to reconstruct 3D scenes from monocular images to generate photorealistic novel perspectives. While neural implicit representations, such as Neural Radiance Fields (NeRF) mildenhall2021nerf , have significantly advanced NVS by modeling radiation fields through multilayer perceptrons (MLPs) and volumetric rendering, these methods struggle to reconcile rendering quality verbin2022ref with computational efficiency fridovich2022plenoxels . Emerging as a novel paradigm, 3D Gaussian Splatting kerbl20233d explicitly parameterizes scenes using adaptive Gaussian primitives. By integrating splat-based rasterization with adaptive density control strategies, 3DGS achieves real-time rendering of unprecedented fidelity while maintaining compact spatial representation.
@@
However, the adaptive density control mechanism in 3DGS exhibits inherent limitations. Specifically, 3DGS’s densification criterion solely relies on the norm of view-space positional gradients, thereby overlooking the directional influence of gradients on the norm computation. During view-space positional gradient calculation for each Gaussian, the superposition of sub-gradient components with consistent directions amplifies the gradient norm. Conversely, counter-directional sub-gradient components mutually attenuate, resulting in reduced gradient norms. These directional conflicts prevent some Gaussians from effective densification, causing over-reconstruction phenomenon and perceptible local blurring. Meanwhile, Gaussians maintaining consistent gradient directions exhibit persistent densification tendencies, causing over-densification phenomena that generate redundant Gaussian primitives and significantly increase memory consumption. Recent work like AbsGS ye2024absgs has proposed using absolute subgradient values to enforce positive gradients for all Gaussians, thereby avoiding subgradient directional conflicts and partially mitigating the over-reconstruction issue. This approach further exacerbates the over-densification phenomenon, as shown in Figure 1.
@@
To address the aforementioned dual challenges, we propose a gradient-direction-aware adaptive density control framework (GDAGS). This approach combines directional subgradient information and gradient magnitude data to establish a novel decision metric. The metric enables controlled densification of Gaussians with conflicting directional gradients while preventing redundant densification in directionally aligned cases. First, GDAGS calculates a normalized GCR to measure the directional consistency of subgradients for each Gaussian. The GCR value approaching 1 indicates highly consistent subgradient directions, whereas a value approaching 0 signifies substantial directional conflict. Subsequently, GDAGS employs GCR to parameterize a nonlinear dynamic weight $w_{i}$, which independently weights the view-space positional gradients of each Gaussian to form a new decision metric controlling densification behavior. During splitting operations, Gaussians exhibiting directional conflicts receive higher weights and are prioritized for splitting, while directionally consistent Gaussians obtain lower weights and are suppressed. Finally, GDAGS introduces a hyperparameter $p$ to explicitly control the number of Gaussians participating in densification, thereby achieving an indirect balance between performance and efficiency. We summarize our contributions as follows:
@@
  • We utilize GCR as a gradient direction consistency measure for adaptive density control to identify Gaussians with consistent and divergent directions.
@@
  • Evaluated on three real-world datasets spanning 13 scenes, our method demonstrated superior visual quality while having a compact spatial representation that requires only 50% Gaussians and memory consumption.
@@ 2.1 Novel view synthesis
Novel View Synthesis has evolved through two interconnected trajectories: advancements in scene representation paradigms and optimizations of rendering efficiency. The field was revolutionized by NeRF mildenhall2021nerf , which introduced MLP-based implicit neural representations to map spatial coordinates to radiance and density values, achieving photorealistic results through volumetric rendering. However, NeRF’s computational demands spurred research bifurcating into quality refinement and efficiency optimization. Approaches like Mip-NeRF barron2021mip enhanced geometric stability using conical ray encoding and distortion regularization, while DVGO sun2022direct and Instant-NGP muller2022instant leveraged explicit 3D grids and hash encoding to achieve near-real-time rendering. This progression culminated in 3D Gaussian Splatting kerbl20233d , which redefined the paradigm by combining explicit Gaussian primitives with differentiable rasterization. By parameterizing scenes as adaptive 3D Gaussians, 3DGS bridged the gap between implicit continuity and explicit computational efficiency, achieving rendering speeds two orders of magnitude faster than NeRF while maintaining visual parity with state-of-the-art methods like Mip-NeRF360 barron2022mip .
@@
The rapid adoption of 3DGS has driven innovations addressing its inherent limitations. Early improvements focused on geometric fidelity: Mip-Splatting yu2024mip integrated multi-scale primitives and low-pass filtering to mitigate aliasing, while GOF yu2024gaussian introduced composite gradient metrics to optimize Gaussian distributions. Subsequent methods explored structural constraints, such as 2DGSs huang20242d planar surface projections and RadeGSs zhang2024rade normal map-guided refinement. Frequency-aware approaches like FreGS zhang2024fregs employed spectral regularization to enhance high-frequency detail recovery, whereas HoGS liuHoGS utilized homogeneous coordinate reparameterization to improve distant object reconstruction. Despite these advances, critical challenges persist in adaptive density control. ABSGS ye2024absgs enforced gradient direction uniformity to resolve over-reconstruction but indiscriminately amplified outliers, while PixelGS zhang2024pixelgs traded memory efficiency for geometric precision by weighting gradients via pixel coverage—a strategy incurring prohibitive storage costs and frame rate penalties. Our work addresses these unresolved trade-offs through directional gradient awareness, systematically balancing precision, efficiency, and memory constraints.
@@ 3 Method
Section 3.1 establishes the methodological motivation for GDAGS, and Section 3.2 outlines 3DGS fundamentals. Sections 3.3.1 and 3.3.2 detail GDAGS’ core components, with full algorithmic implementation provided in Appendix A. The complete pipeline is shown in Figure 2.
Figure 2: The pipeline of GDAGS. First, for each Gaussian, GDAGS computes the GCR to quantify the directional coherence of its subgradients. Subsequently, this GCR metric is mapped through a nonlinear dynamic weighting function to generate per-Gaussian gradient weights, which modulate the view-space positional gradient magnitudes and produce a refined decision metric. Finally, this decision metric is compared against a predefined threshold to dynamically regulate densification. The detailed algorithm pseudocode is presented in Appendix A.
3.1 Motivation
The density control mechanism in 3D Gaussian Splatting fundamentally relies on comparing view-space positional gradient norms with predefined thresholds. However, the directional characteristics of subgradients implicitly modulate these gradient norms, either attenuating or amplifying the magnitude. This manifests as undesirable phenomena, including local blurring and excessive Gaussian proliferation, ultimately degrading rendering quality. To address these limitations, we propose to explicitly incorporate subgradient directional information into a nonlinear weighting function to construct a novel decision metric for resolving over-reconstruction and over-densification issues. Specifically, during splitting operations (targeting larger Gaussians), the GCR is utilized to determine the weighting strategy: Gaussians with smaller GCR values (indicating directional conflicts) receive amplified decision metrics through nonlinear weighting, thereby prioritizing their densification. Conversely, Gaussians with larger GCR values (exhibiting directional consistency) obtain diminished decision metrics and are suppressed. In contrast, during cloning operations (targeting smaller Gaussians), the weighting mechanism is inverted: low-GCR Gaussians are suppressed while high-GCR Gaussians are promoted for densification. This is because the Gaussian clones of high GCR can smoothly move towards the gradient direction, while the Gaussian clones of low-GCR stack locally. By systematically integrating both gradient directional information and magnitude characteristics into the decision metric, GDAGS optimizes density control mechanisms. This approach not only mitigates the inherent limitations of conventional 3DGS but also significantly reduces both the required number of Gaussians and memory consumption. Additional technical details are provided in Appendix B.
@@ 3.2 Preliminary
3DGS represents scenes using a set of learnable anisotropic 3D Gaussian primitives. Each Gaussian is parameterized by a position $\mu\in\mathbb{R}^{3}$, an opacity $\alpha\in[0,1]$, a covariance matrix $\Sigma\in\mathbb{R}^{3\times 3}$, and spherical harmonics (SH) arfken2011mathematical coefficients for view-dependent color. The covariance matrix $\Sigma$ is decomposed into a rotation matrix $R$ and a scaling matrix $S$ to ensure positive semi-definiteness:
@@
$C(\mathbf{x}^{\prime})=\sum_{k=1}^{K}c_{k}\alpha_{k}G^{\prime}k(\mathbf{x}^{\prime})T_{k},\quad T_{k}=\prod{j=1}^{k-1}\left(1-\alpha_{j}G^{\prime}_{j}(\mathbf{x}^{\prime})\right),$ (3)
@@
To address sparse input from structure-from-motion (SfM) schonberger2016structure , a densification strategy is employed. Under-reconstructed regions (missing geometric features) are addressed by cloning Gaussians, increasing their number and volume. Over-reconstructed regions (large Gaussians covering small areas) are split into smaller Gaussians, maintaining volume but increasing count. Densification decisions are based on view-space gradients. For a Gaussian $G_{i}$ with projection $\mu_{i}^{k}$ under viewpoint $k$ and loss $L_{k}$, the average gradient magnitude is:
@@
where $M$ is the number of viewpoints involved in training. Split Gaussians into smaller components if $\nabla_{\mu_{i}}L>\tau_{p}$ and $\Sigma_{3D}^{i}>\tau_{s}$. Clone Gaussians to preserve local density if $\nabla_{\mu_{i}}L>\tau_{p}$ and $\Sigma_{3D}^{i}\leq\tau_{s}$.
@@ 3.3.1 Directional consistency measurement
The denominator $\sum_{pixel}\|\nabla_{i,pixel}^{v}\|_{2}$ captures the total gradient activity agnostic to directional variations, while the numerator $\|\sum_{pixel}\nabla_{i,pixel}^{v}\|_{2}$ quantifies the net directional alignment.
@@ 3.3.2 Nonlinear dynamic weight design
Let $\theta_{i,v}\in[0,2\pi)$ denote the gradient direction of Gaussian $i$ in view $v$, modeled as samples from a von Mises-Fisher (vMF) distribution:
$p(\theta_{i,v}|m_{i},\kappa_{i})=\frac{\kappa_{i}}{2\pi I_{0}(\kappa_{i})}e^{\kappa_{i}\cos(\theta_{i,v}-m_{i})},$ (6)
where $m_{i}$ is the mean direction and $\kappa_{i}>0$ the concentration parameter. The empirical directional consistency $\mathcal{C}_{i}$ relates to $\kappa_{i}$ through:
$\mathcal{C}_{i}=\frac{I_{1}(\kappa_{i})}{I_{0}(\kappa_{i})}\approx 1-\frac{1}{2\kappa_{i}}+\mathcal{O}(\kappa_{i}^{-2}),$ (7)
where $I_{n}$ denotes the modified Bessel function of order $n$.
Then, we define the target distribution $q^{*}(\nabla)$ that emphasizes directionally inconsistent regions:
$q^{*}(\nabla)=\frac{1}{Z}\|\nabla\|_{2}^{1-\mathcal{C}_{i}}\quad\text{with}\quad Z=\int\|\nabla\|_{2}^{1-\mathcal{C}_{i}}d\nabla,$ (8)
where $Z$ is the normalization constant. The KL divergence between empirical distribution $p(\nabla)$ and $q^{*}(\nabla)$ is:
$D_{KL}(p\|q^{*})=\mathbb{E}_{p}[\log p(\nabla)]-\mathbb{E}_{p}[\log\|\nabla\|_{2}^{1-\mathcal{C}_{i}}]+\log Z,$ (9)
Substituting the weighted distribution $p_{w}(\nabla)\propto w(\nabla)p_{0}(\nabla)$ and taking variational derivative:
$\frac{\delta D_{KL}}{\delta w}=\frac{(1-\mathcal{C}_{i})\log\|\nabla\|_{2}-\log w}{\mathbb{E}[w]}+\lambda=0,$ (10)
where $\lambda$ is a Lagrange multiplier. Solving this yields the optimal weight:
$w^{*}(\nabla)=\exp\left((1-\mathcal{C}_{i})\log\|\nabla\|_{2}\right).$ (11)
For ease of use and simplification of calculations, we represent $w^{*}(\nabla)$ as $w_{i}\propto(1-\mathcal{C}_{i})$, where $w_{i}$ is the weight corresponding to Gaussian $G_{i}$. Furthermore, we design $w_{i}$ as follows:
@@
where $\alpha\in[0,1]$ acts as a baseline suppress factor that suppresses split operations under directional consistency, while $\beta$ serves as an amplification factor to enhance weight responses when directional discrepancies occur. The shape-adjustable hyperparameter $p$ governs the nonlinearity of the weight response curve. As $p$ increases, the curve exhibits progressively steeper gradients, which correspondingly diminish the proportion of Gaussians surpassing the split threshold. This formulation preserves three fundamental mathematical properties: monotonicity, boundedness, and convexity. For visual reference, the response curves corresponding to different $p$ values are presented in Appendix B. The decision measurements are represented as follows:
$\nabla_{split}=\nabla_{\mu_{i}}L\cdot w_{i},\nabla_{clone}=\nabla_{\mu_{i}}L/w_{i}.$ (13)

### v2

@@ Abstract
The emergence of 3D Gaussian Splatting (3DGS) has significantly advanced Novel View Synthesis (NVS) through explicit scene representation, enabling real-time photorealistic rendering. However, existing approaches manifest two critical limitations in complex scenarios: (1) Over-reconstruction occurs when persistent large Gaussians cannot meet adaptive splitting thresholds during density control. This is exacerbated by conflicting gradient directions that prevent effective splitting of these Gaussians; (2) Over-densification of Gaussians occurs in regions with aligned gradient aggregation, leading to redundant component proliferation. This redundancy significantly increases memory overhead due to unnecessary data retention. We present Gradient-Direction-Aware Gaussian Splatting (GDAGS) to address these challenges. Our key innovations: the Gradient Coherence Ratio (GCR), computed through normalized gradient vector norms, which explicitly discriminates Gaussians with concordant versus conflicting gradient directions; and a nonlinear dynamic weighting mechanism leverages the GCR to enable gradient-direction-aware density control. Specifically, GDAGS prioritizes conflicting-gradient Gaussians during splitting operations to enhance geometric details while suppressing redundant concordant-direction Gaussians. Conversely, in cloning processes, GDAGS promotes concordant-direction Gaussian densification for structural completion while preventing conflicting-direction Gaussian overpopulation. Comprehensive evaluations across diverse real-world benchmarks demonstrate that GDAGS achieves superior rendering quality while effectively mitigating over-reconstruction, suppressing over-densification, and constructing compact scene representations.
@@ 1 Introduction
Figure 1: (a) illustrates the Gaussian ellipsoid splatting process, where arrows of different colors represent the gradient direction and magnitude of different Gaussians on the pixels. (b) shows the densification process of different methods. In 3DGS, a large Gaussian covering many pixels may fail to split because the combined gradient magnitude from different pixels falls below the threshold, leading to over-reconstruction as shown in the Rendered part of (c), which manifests as blurry areas. AbsGS forces all Gaussian gradients to be positive, causing the combined gradient magnitude from different pixels to increase significantly. This results in a substantial rise in the number of splitting Gaussians, leading to over-densification as shown in the Top view of the point cloud part of (c), which appears as a large number of Gaussian point clouds outside the scene. Our proposed method utilizes gradient direction information to effectively control the splitting of large Gaussians and the growth in the number of small and medium Gaussians. (c) also presents the performance (performance metrics are labeled in the top-left corner) and efficiency (FPS and total number of Gaussians in the scene, #G) of different methods.
@@
Novel View Synthesis (NVS) constitutes a foundational challenge in computer graphics and vision, aiming to reconstruct 3D scenes from monocular images to generate photorealistic novel perspectives. While neural implicit representations, such as Neural Radiance Fields (NeRF) (Mildenhall et al., 2021), have significantly advanced NVS by modeling radiation fields through multilayer perceptrons (MLPs) and volumetric rendering, these methods struggle to reconcile rendering quality (Verbin et al., 2022) with computational efficiency (Fridovich-Keil et al., 2022). Emerging as a novel paradigm, 3D Gaussian Splatting (3DGS) (Kerbl et al., 2023) explicitly parameterizes scenes using adaptive Gaussian primitives. By integrating splat-based rasterization with adaptive density control strategies, 3DGS achieves real-time rendering of unprecedented fidelity while maintaining compact spatial representation.
@@
However, the adaptive density control mechanism in 3DGS exhibits inherent limitations (Ye et al., 2024; Huang et al., 2025; Zeng et al., 2025). Specifically, 3DGS’s densification criterion solely relies on the norm of view-space positional gradients, thereby overlooking the directional influence of gradients on the norm computation. During view-space positional gradient calculation for each Gaussian, the superposition of sub-gradient components with consistent directions amplifies the gradient norm. Conversely, counter-directional sub-gradient components mutually attenuate, resulting in reduced gradient norms. These directional conflicts prevent some Gaussians from effective densification, causing the over-reconstruction phenomenon and perceptible local blurring (Zhang et al., 2024b; Zhou et al., 2025). Meanwhile, Gaussians maintaining consistent gradient directions exhibit persistent densification tendencies, causing over-densification phenomena that generate redundant Gaussian primitives and significantly increase memory consumption. Recent work like AbsGS (Ye et al., 2024) has proposed using absolute subgradient values to enforce positive gradients for all Gaussians, thereby avoiding subgradient directional conflicts and partially mitigating the over-reconstruction issue, but this approach further exacerbates the over-densification phenomenon.
@@
To address the aforementioned dual challenges, we propose a gradient-direction-aware adaptive density control framework (GDAGS). This approach combines directional gradient information and gradient magnitude data to establish a novel decision metric. The metric enables controlled densification of Gaussians with conflicting directional gradients while preventing redundant densification in directionally aligned cases. First, GDAGS calculates GCR to measure the directional consistency of subgradients for each Gaussian. The GCR approaching 1 indicates highly consistent subgradient directions, whereas a value approaching 0 signifies substantial directional conflict. Subsequently, GDAGS employs GCR to parameterize a nonlinear dynamic weight $w_{i}$, which independently weights the view-space positional gradients of each Gaussian to form a new decision metric controlling densification behavior. During splitting operations, Gaussians exhibiting directional conflicts receive higher weights and are prioritized for splitting, while directionally consistent Gaussians obtain lower weights and are suppressed. Finally, GDAGS introduces a hyperparameter $p$ to explicitly control the number of Gaussians participating in densification, thereby achieving an indirect balance between performance and efficiency. Comparative results and efficiency analyses are summarized in Figure 1. We summarize our contributions as follows:
@@
  • We propose GCR as a gradient direction consistency measure for adaptive density control to identify Gaussians with consistent and divergent directions.
@@
  • Evaluated on three real-world datasets spanning 13 scenes, our method demonstrates superior visual quality while having a compact spatial representation.
@@ 2.1 Novel view synthesis
Novel View Synthesis has evolved through two interconnected trajectories: advancements in scene representation paradigms and optimizations of rendering efficiency. The field was revolutionized by NeRF (Mildenhall et al., 2021), which introduced MLP-based implicit neural representations to map spatial coordinates to radiance and density values, achieving photorealistic results through volumetric rendering. However, NeRF’s computational demands spurred research bifurcating into quality refinement and efficiency optimization. Approaches like Mip-NeRF (Barron et al., 2021) enhanced geometric stability using conical ray encoding and distortion regularization, while DVGO (Sun et al., 2022) and Instant-NGP (Müller et al., 2022) leveraged explicit 3D grids and hash encoding to achieve near-real-time rendering. This progression culminates in 3D Gaussian Splatting (Kerbl et al., 2023), which redefined the paradigm by combining explicit Gaussian primitives with differentiable rasterization. By parameterizing scenes as adaptive 3D Gaussians, 3DGS bridged the gap between implicit continuity and explicit computational efficiency, achieving rendering speeds two orders of magnitude faster than NeRF while maintaining visual parity with state-of-the-art methods like Mip-NeRF360 (Barron et al., 2022).
@@
The rapid adoption of 3DGS has driven innovations addressing its inherent limitations (Cheng et al., 2024; Hamdi et al., 2024; Kheradmand et al., 2024; Lee et al., 2024; Li et al., 2025b). Early improvements focus on geometric fidelity: Mip-Splatting (Yu et al., 2024a) integrates multi-scale primitives and low-pass filtering to mitigate aliasing, while GOF (Yu et al., 2024b) introduces composite gradient metrics to optimize Gaussian distributions. Subsequent methods explore structural constraints, such as 2DGSs (Huang et al., 2024) planar surface projections and RadeGSs (Zhang et al., 2024a) normal map-guided refinement. Frequency-aware approaches like FreGS (Zhang et al., 2024b) employ spectral regularization to enhance high-frequency detail recovery, whereas HoGS (Liu et al., 2025) utilizes homogeneous coordinate reparameterization to improve distant object reconstruction. Despite these advances, critical challenges persist in adaptive density control. AbsGS (Ye et al., 2024) enforces gradient direction uniformity to resolve over-reconstruction but indiscriminately amplifies outliers, while PixelGS (Zhang et al., 2024c) trades memory efficiency for geometric precision by weighting gradients via pixel coverage—a strategy incurring prohibitive storage costs and frame rate penalties. ReAct-GS (Cheng et al., 2025) proposes a densification and reactivation mechanism based on importance perception to improve rendering quality, while PSRGS (Li et al., 2025a) solves the problem of high-frequency detail loss by identifying geometric and texture coupling regions. These methods inevitably increase model complexity by introducing additional modules or strategies. Our work addresses these unresolved trade-offs through directional gradient awareness, systematically balancing precision, efficiency, and memory constraints.
@@ 3 Method
Section 3.1 establishes the methodological motivation for GDAGS, and Section 3.2 outlines 3DGS fundamentals. Sections 3.3 detail GDAGS’ core components, with full algorithmic implementation provided in Appendix B. The complete pipeline is shown in Figure 2.
Figure 2: The pipeline of GDAGS. First, for each Gaussian, GDAGS computes the GCR to quantify the directional coherence of its subgradients. Subsequently, this GCR metric is mapped through a nonlinear dynamic weighting function to generate per-Gaussian gradient weights, which modulate the view-space positional gradient magnitudes and produce a refined decision metric. Finally, this decision metric is compared against a predefined threshold to dynamically regulate densification.
@@ 3.1 Preliminary
3DGS represents scenes using a set of learnable anisotropic 3D Gaussian primitives. Each Gaussian is parameterized by a position $\mu\in\mathbb{R}^{3}$, an opacity $\alpha\in[0,1]$, a covariance matrix $\Sigma\in\mathbb{R}^{3\times 3}$, and Spherical Harmonics (SH) (Arfken et al., 2011) coefficients for view-dependent color. The covariance matrix $\Sigma$ is decomposed into a rotation matrix $R$ and a scaling matrix $S$ to ensure positive semi-definiteness:
@@
$C(\mathbf{x}^{\prime})=\sum_{k=1}^{K}c_{k}\alpha_{k}G^{\prime}k(\mathbf{x}^{\prime})T_{k},\quad T_{k}=\textstyle\prod_{j=1}^{k-1}\left(1-\alpha_{j}G^{\prime}_{j}(\mathbf{x}^{\prime})\right),$ (3)
@@
To address sparse input from structure-from-motion (SfM) (Schonberger and Frahm, 2016), a densification strategy is employed. Under-reconstructed regions (missing geometric features) are addressed by cloning Gaussians, increasing their number and volume. Over-reconstructed regions (large Gaussians covering small areas) are split into smaller Gaussians, maintaining volume but increasing count. Densification decisions are based on view-space gradients. For a Gaussian $G_{i}$ with projection $\mu_{i}^{k}$ under viewpoint $k$ and loss $L_{k}$, the average gradient magnitude is:
@@
where $M$ is the number of viewpoints involved in training.
Split Gaussians into smaller components if $\nabla_{\mu_{i}}L>\tau_{p}$ and $\Sigma_{3D}^{i}>\tau_{s}$. Clone Gaussians to preserve local density if $\nabla_{\mu_{i}}L>\tau_{p}$ and $\Sigma_{3D}^{i}\leq\tau_{s}$.
3.2 Motivation
The density control in 3D Gaussian Splatting primarily depends on comparing the norms of view-space positional gradients against predefined thresholds. However, the directional properties of these gradients implicitly influence their magnitudes—either attenuating or amplifying them—leading to undesirable effects such as local blurring and excessive Gaussian proliferation, which ultimately impair rendering quality. To overcome these issues, we introduce a novel decision metric that explicitly incorporates subgradient direction through a nonlinear weighting function, effectively addressing both over-reconstruction and over-densification. Specifically, during splitting operations—which target larger Gaussians—we use the GCR to guide the weighting strategy. Gaussians with smaller GCR values, indicating directional conflicts, receive an amplified decision metric via nonlinear weighting, thus prioritizing their densification. In contrast, those with larger GCR values, reflecting directional consistency, are assigned a diminished metric and suppressed. During cloning operations, which affect smaller Gaussians, the logic is reversed: low-GCR Gaussians are suppressed, while high-GCR Gaussians are promoted. This is because high-GCR clones can smoothly propagate along the gradient direction, whereas low-GCR clones tend to accumulate locally without meaningful displacement. By systematically integrating both gradient direction and magnitude into the densification criterion, GDAGS optimizes the density control mechanism. This approach not only mitigates key limitations of standard 3DGS but also significantly reduces the number of Gaussians required and corresponding memory usage.
@@ 3.3.1 Directional consistency measurement
The denominator $\sum_{pixel}\|\nabla_{i,pixel}^{v}\|_{2}$ captures the total gradient activity agnostic to directional variations, while the numerator $\|\sum_{pixel}\nabla_{i,pixel}^{v}\|_{2}$ quantifies the net directional alignment. Due to the fact that GCR calculates subgradients for each pixel in multiple views, GDAGS works under a clear assumption that most Gaussian distributions in the scene can receive sufficient gradient information.
@@ 3.3.2 Nonlinear dynamic weight design
The core of our adaptive density control lies in transforming the GCR into a dynamic weight that modulates the gradient magnitude used in densification decisions. A naive linear weighting scheme proves insufficient, as it lacks the expressive power to aggressively suppress high-consistency Gaussians while sensitively amplifying those with critical directional conflicts. To address this, we introduce a nonlinear weighting function designed with two key objectives: strongly attenuating the contribution of Gaussians with high GCR values to prevent redundant densification in well-defined regions. And strategically amplify the influence of Gaussians with intermediate to low GCR values, prioritizing them for densification to resolve directional conflicts and enhance geometric detail. We design a nonlinear weighting function defined as:
@@
where $\mathcal{C}_{i}$ is the gradient coherence ratio for Gaussian $i$ computed via Equation 5, $\alpha$ is the fundamental inhibitory factor that suppresses Gaussian densification with consistent gradient directions during the splitting process, and suppresses Gaussian densification with inconsistent gradient directions during the cloning process. $\beta$ acts as the amplification factor, scaling the weight for Gaussians with strong directional conflicts, and $p$ governs the steepness of the exponential suppression, rapidly decreasing the weight for Gaussians with high directional consistency.
The exponential form is motivated by the need for a strong, monotonic decay that effectively neutralizes the influence of Gaussians whose gradients are already aligned. This prevents the unnecessary splitting or cloning that leads to over-densification. The power-law form in the amplification region, $(1-\mathcal{C}_{i})^{p}$, is chosen because it is more sensitive to changes in $\mathcal{C}_{i}$ when consistency is low compared to a linear function. This sensitivity is crucial for identifying and prioritizing Gaussians that are large and cover areas with complex geometry (and hence conflicting gradients), which are most in need of splitting.
The modulated gradient norm for Gaussian $i$ is then computed as:
$\tilde{\nabla}_{\mu_{i}}L=w_{i}\cdot\nabla_{\mu_{i}}L.$ (7)
This refined metric replaces the original gradient norm in all densification decisions, enabling consistent gradient-direction-aware control during optimization. The weighting function is applied strategically according to the densification operation: during splitting, the weight $w_{i}$ is applied directly to prioritize the fragmentation of large Gaussians exhibiting directional conflicts. In cloning operations, the inverse policy $(1/w_{i})$ is adopted to encourage the propagation of small Gaussians with high GCR along surface directions, while suppressing those with low GCR to prevent chaotic local accumulation. Comprehensive ablation studies in Section 4.3.1 confirm the superiority of this nonlinear weighting strategy over linear alternatives, and a sensitivity analysis of parameters $p$ and $\beta$ is provided in Section 4.3.2.
4 Experiments
4.1 Setup
Datasets. Similar to the 3DGS dataset, we select all 9 scenes introduced in MIP-NERF360 (Barron et al., 2022), including indoor and outdoor scenes, Train and Truck scenes in Tanks&Temples (Hedman et al., 2018), and drJohnson and playroom scenes in Deep Blending (Knapitsch et al., 2017).
Baselines. The proposed method is benchmarked against three categories of approaches: fast NeRF frameworks (Plenoxels (Fridovich-Keil et al., 2022), Instant-NGP (Müller et al., 2022)), state-of-the-art NeRF-based models (Mip-NeRF360 (Barron et al., 2022)), and 3DGS variants including the widely adopted 3DGS (Kerbl et al., 2023), Taming 3DGS (Mallick et al., 2024) and mini-splatting (Fang and Wang, 2024), over-reconstruction-focused AbsGS (Ye et al., 2024), and Pixel-GS (Zhang et al., 2024c), ensuring comprehensive evaluation across all datasets.
Implementation Details. Our experimental setup follows the original paper parameter settings of 3DGS, with densification performed every 100 iterations, stopping densification after 15k iterations and completing training at 30k iterations. The suppression factor $\alpha$ in the weight function is set to 0.8, the amplification factor $\beta$ is set to the size threshold of 25, and the hyperparameter $p$ is set to 15. All experiments are conducted on a single NVIDIA 4090 GPU with 24GB of memory. We report the view metrics PSNR, SSIM, and LPIPS (Zhang et al., 2018), and the memory used to store Gaussian parameters to demonstrate the trade-off between performance and efficiency.